# Applying statistical learning theory to deep learning

### v1

@@ 4.1 Mirror Descent
We derived mirror descent as regularizing this object with respect to optimizing the first order object. We wrote it in the following form:
$w_{k+1}$ $=\arg\min_{w\in\mathcal{W}}\langle\nabla f(w_{k},z_{k}),w\rangle+\lambda_{k}D_{\Psi}(w||w_{k})$ (45)
$=\Pi_{\Psi}^{{\mathcal{W}}}\left(\nabla\Psi^{-1}\left(\nabla\Psi(w_{t})-\frac{1}{\lambda_{t}}\nabla f(w_{t},z_{t})\right)\right)$ (46)
@@
$D_{\Psi}(w||w^{\prime})=\Psi(w)-(\Psi(w^{\prime})+\langle\nabla\Psi(w^{\prime}),w-w^{\prime}\rangle)$
we replace $\lambda_{k}$ with $\eta_{k}$ since the regularization is linked to the step size, as was discussed in the previous lecture.
@@
We cannot directly add points in the primal and dual spaces, so we use a link function. To go from primal space to dual space, we use the gradient of the potential, and to go back to the primal space from the dual space,
@@
We can write the gradient descent problem approximately as:
$w_{k+1}$ $=\arg\min_{w\in\mathcal{W}}\langle\nabla f(w_{k},z_{k}),w\rangle+\frac{1}{\eta_{k}}(w-w_{k})^{T}\nabla^{2}\Psi(w_{k})(w-w_{k})$ (47)
$=w_{k}-\eta_{k}\nabla^{2}\Psi(w_{k})^{-1}\nabla f(w_{k},z_{k})\,,\quad\text{define}\,\rho(w_{k})\coloneqq\nabla^{2}\Psi(w_{k})^{-1}$ (48)
The latter version of the update rule is known as natural gradient descent, which was first popularized in the context of information geometry by S. Amari (see [1]). It is valid for any potential $\Psi$ (or really for any metric tensor). For intuition, consider that there is a manifold with some local metric $\rho\,,$ which is the Hessian of $\Psi\,,$ and we optimize on the manifold. Thus, here we take the view that natural gradient descent is an approximation to mirror descent.
@@
As the step size goes to 0, the stochasticity goes away: in any time interval, as the step size goes to 0, we take more steps and so we see more samples, so the stochasticity goes away. Thus, instead of writing the stochastic gradient, we can write the population gradient:
@@
There are two ways to interpret what is hapening here: when the step size goes to 0, both mirror descent and natural gradient descent (which are the same in the infinitesimal step size limit) converge to the Riemannian gradient flow on the population objective. (Note: this is another way to think about what we are doing with SGD, because as the step size gets smaller and smaller, we are approaching optimizing the population loss.) The other way to think about natural gradient descent and mirror descent is: we endow the space with some local geometry and take steps based on that local geometry. Natural gradient descent corresponds to a forward Euler discretization of the natural gradient flow rule (Eqn. 49):
@@ 4.1.1 Examples of Mirror Descent
which holds as long as $\Psi$ is 1-strongly convex with respect to our choice of norm. Here, $w^{\star}$ is no longer the best in the norm bounded class, since we do not wish to limit the norm explicitly. The guarantee above means that we can compete with any $w^{\star}$, we just have to pay for its complexity in the second term.
Question
When the metric tensor doesn’t correspond to a Hessian, does that mean we cannot come up with a global complexity measure? Answer The answer is a combination of “no” and “I don’t know .”
@@ 4.1.2 Smoothness and Batching
$\mathbb{E}\left[F(\bar{w}_{k})\right]\leq F(w^{\star})+O\left(\frac{H\Psi^{\star}(w)}{k}+\sqrt{\frac{\sigma^{2}\Psi^{\star}(w)}{k}}\right)\,,\text{where }\sigma\text{ s.t. }\mathbb{E}\left[\left|\left|\nabla f(w,z)-\nabla F(z)\right|\right|^{2}\right]\leq\sigma^{2}$ (55)
@@
We can see here is that optimization is easier than the statistical aspect. It never really helps us to take more steps than the number of samples. In the non-convex case, this is substantially different: in particular, the optimization might be harder than the statistical part, and so it might be worth viewing the samples multiple times.
@@
The guarantee in Eqn. 55 holds when $H$ is the relative smoothness parameter.
@@ 4.2 General Steepest Descent
So far, we have talked about mirror descent and have related it to Riemannian gradient flow. We can also think more generally about steepest descent methods that are not related to a metric.
@@ 4.3 Implicit bias of descent methods
So far, we have discussed the convex setting, and we related optimization guarantees and generalization guarantees. We get generalization guarantees with respect to $w^{\star}$ considering only the dynamics of optimization. We studied the connection between the geometry of searching the space and the corresponding complexity. However, there are some limitations to what we have seen so far. First of all, this is only the convex case. Second of all, this isn’t what we’re seeing in deep learning. Here, we are not driving training error to 0. The generalization seems to rely on the fact that we are using stochasticity. That is different from what we see in the examples from the previous lectures where we train with full batch, multi-pass gradient methods but still get good generalization.
@@
$w_{\lambda}=\arg\min\hat{F}_{k}(w)+\lambda\Psi(w)\,.$
And secondly, consider $\bar{w}_{k}$, the solution achieved after $k$ iterations of mirror descent. In either of these cases, we get the same generalization guarantee, but for different solutions. In the Lipschitz case with optimally chosen $\lambda$, the distance (suboptimality, really) between $w_{\lambda}$ and $w^{\star}$ is $1/\sqrt{k}\,.$ The distance between $w^{\star}$ and $\bar{w}_{k}$ is also $1/\sqrt{k}\,,$ but the distance between $w_{\lambda},\bar{w}$ is also $1/\sqrt{k}$. We know this because we know that we can view one-pass of this as optimizing the training objective, which in $k$ steps gets suboptimality $1/\sqrt{k}\,.$ Thus, we are not saying that implicit regularization gets us to the same solution as explicit regularization, but rather we are saying that the generalization guarantees hold. If we keep repeating passes, we might get to minimizer of the training error, but it’s unclear if this is beneficial.
@@ 4.3.1 Deriving Implicit Regularization for Gradient Descent
We’ve dropped $\phi$ for ease of notation, but we should think of $x$ as being some feature map. Let $w\in\mathbb{R}^{d}\,,d>>m\,.$ Let us use gradient descent:
$w_{k+1}=w_{k}-\eta\nabla\hat{F}(w_{k})\,;w_{0}=0\,.$
We know from previous discussion that this will result in implicit $\ell_{2}$ regularization, but we will formally derive this. Additionally, this will later allow us to understand what happens when using mirror descent.
@@
This is because each update involves adding a scaled gradient, and the gradient is dependent on a residual. This already tells us a lot about what we will get: even though we have $d$ parameters, we only will ever be in an $m$-dimensional subspace. We will also assume that we get to a global optimum. This is not always easy to show but are going to assume it. Since we are in the overparametrized setting, we know that $Xw=y\,,$ where $w$ is the predictor to which we finally converge. We claim that the point we get is exactly the minimizer of the following optimization problem:
@@
We have the optimization that we are presumably optimizing: that is, minimizing the empirical error, and we have the optimization problem above. In general, these arguments are going to proceed by claiming that solving the first optimization problem actually implicitly solves the second optimization problem. To show this, we use the method of Karush-Kuhn-Tucker (KKT) conditions for optimality of a solution to a constrained optimization problem. The optimum for a constrained optimization problem is uniquely characterized by its KKT conditions. Let us look at the KKT conditions. We introduce dual variables $\nu$ for the constraints, giving us that the Lagrangian is:
@@
We call $\nu$ the dual variables, or the variables in the dual space. We observe that the stationarity condition shows exactly what we claimed earlier, that $w_{k}$ are in the span of the data, and the primal feasibility condition finds the interpolator. Since a point that satisfies these two conditions is optimal for the optimization problem in Eqn. 57, and since gradient descent that arrives at a 0 training error predictor satisfies these two conditions, gradient descent finds an optimal solution to the problem in Eqn. 57, i.e., a minimum Euclidean norm solution.
@@ 4.3.2 Similar Argument for Mirror Descent
With this argument in mind, we can apply a similar procedure to analyze the output of mirror descent. We are going to show that with mirror descent, we get to a solution that is good with respect to the stated potential. To show this, we will use essentially the same proof. The optimization problem is still the same as before: we are minimizing the training objective, but the optimization algorithm will differ. Now, we optimize $\hat{F}$ using mirror descent with respect to some potential function $\Psi$. Let us reproduce the iterates:
@@
$w_{k+1}=\nabla\Psi^{-1}\bigg{(}\nabla\Psi(w_{0})-\sum_{i=1}^{k}\nabla\hat{F}(w_{k})\bigg{)}$
That is, the $k+1$ iterate is the mapping back into the primal space of the accumulation of the gradients in the dual space. As before, we can see that the gradients lie in the span of the data (in the dual space). Thus, again, $w_{k}\in\mathcal{M}=\{\nabla\Psi^{-1}(\nabla\Psi(w_{0})+X^{T}s)\,\forall\,x\in\mathbb{R}^{m}\}\,.$ This is not a flat manifold (zero curvature) in the primal space, but it is in the dual space. To see what point we converge to, we use the fact that in the end, we converge to a global optimum. This imposes $m$ linear constraints, which when intersected with an $m$-dimensional manifold gives us a unique point. To determine which unique point that is, we write an optimization problem whose KKT conditions match the two sets of constraints (global optimality and lying in the manifold):
@@
4.3.3 General Method
We started from the optimization problem of minimizing the training objective and then saw trajectory stays within a manifold. The second set of constraints we imposed came from the global optimality of the solution reached. We then matched these two sets of constraints to KKT conditions for some other optimization problem.
If $w_{0}$ were the minimizer of the potential function, then the optimization problem becomes minimizing the potential function subject to $Xw=y$.
@@ 5 Lecture 5: Implicit bias with linear functionals and the square loss
The goal of this lecture is to provide analytical evidence to the lazy [8] and rich regimes in learning problems with gradient descent. We will consider a simple model for which the dynamics of gradient descent may be solved exactly and we will study these trajectories as a function of the magnitude of the initialization and model architectures, leading to various implicit biases.
@@ 5.1 Setting
We consider models parametrized by a weight vector $\mathbf{w}\in\mathbb{R}^{p}$ acting on an input space $\mathcal{X}$ and denote $F(\mathbf{w})\in\{f:\mathcal{X}\to\mathbb{R}\}$ the predictor implemented by $\mathbf{w}$, such that $F(\mathbf{w})(\mathbf{x})=f(\mathbf{w},\mathbf{x})$. We will focus on models linear in $\mathbf{x}$ represented by a linear functional in the dual space of $\mathcal{X}$, denoted $\mathcal{X}^{*}$ represented by a vector $\boldsymbol{\beta}_{\mathbf{w}}$ such that $f(\mathbf{w},\mathbf{x})=\langle\boldsymbol{\beta}_{\mathbf{w}},\mathbf{x}\rangle$, i.e. $\boldsymbol{\beta}_{\mathbf{w}}$ is some transform, potentially non-linear of $\mathbf{w}$. We consider the supervised learning problem with $n$ sample pairs $(\mathbf{x}_{i},y_{i})$ where $y_{i}\in\mathbb{R}$ is a response vector. The parameters are learned by minimizing the empirical loss
@@
We assume the model is homogeneous of order $D$, i.e., for any constant $c>0$ $F(c\mathbf{w})=C^{D}h_{\mathbf{w}}$. The order $D$ is related to the depth of networks with homogeneous activations (e.g. a linear or Relu). A linear model is homogeneous of order $1$, a factorization model of order $2$. Our goal is three-fold :
  • first, we want to characterize the implicit bias of gradient descent in this setup. In regards to the previous lectures, is optimzing in the parameters space using gradient descent equivalent to optimizing in the functional space w.r.t. some metric tensor and potential ? What is this potential
  • secondly is it equivalent to explicit regularization ? For instance, we saw that GD on a least-squares problem converges towards the minimum $\ell_{2}$ norm solution, equivalent to explicitly penalizing the norm. Same here ?
  • finally, we would like to study the transition between kernel (a.k.a. lazy) regime and rich regime. What parameters govern this behaviour ? We have seen in the previous lectures that in many cases we obtain implicit biases that cannot be reached with kernel methods (sparsity, nuclear norm, …). Can we get a more precise picture on simple models ?

### v2

@@ 4.1 Mirror Descent
We derived mirror descent as iteratively minimizing a regularized form of the first order approximation of the objective function using a convex potential. We obtained the following form:
$w_{k+1}$ $=\arg\min_{w\in\mathcal{W}}\langle\nabla f(w_{k},z_{k}),w\rangle+\lambda_{k}D_{\Psi}(w||w_{k})$
$=\Pi_{\Psi}^{{\mathcal{W}}}\left(\nabla\Psi^{-1}\left(\nabla\Psi(w_{k})-\frac{1}{\lambda_{k}}\nabla f(w_{k},z_{k})\right)\right)$ (45)
@@
$D_{\Psi}(w||w^{\prime})=\Psi(w)-\Psi(w^{\prime})+\langle\nabla\Psi(w^{\prime}),w-w^{\prime}\rangle.$
Since the value of the step size $\eta_{k}$ is directly linked to the value of the regularization parammeter $\lambda_{k}$, we may rewrite the iteration as
$w_{k+1}=\Pi_{\Psi}^{{\mathcal{W}}}\left(\nabla\Psi^{-1}\left(\nabla\Psi(w_{k})-\eta_{k}\nabla f(w_{k},z_{k})\right)\right).$ (46)
@@
We can then write the mirror descent problem approximately as:
$w_{k+1}$ $=\arg\min_{w\in\mathcal{W}}\langle\nabla f(w_{k},z_{k}),w\rangle+\frac{1}{\eta_{k}}(w-w_{k})^{T}\nabla^{2}\Psi(w_{k})(w-w_{k})$
$=w_{k}-\eta_{k}\nabla^{2}\Psi(w_{k})^{-1}\nabla f(w_{k},z_{k})\,,$
$\text{where}\quad\rho(w_{k})\coloneqq\nabla^{2}\Psi(w_{k})^{-1}$
The latter version of the update rule is known as natural gradient descent, which was first popularized in the context of information geometry by S. Amari (see [1]). It is valid for any sufficiently regular potential $\Psi$ (or really for any metric tensor). For intuition, consider that there is a manifold with some local metric $\rho\,,$ which is the Hessian of $\Psi\,,$ and we optimize on this manifold. Thus, here we adopt the point of view that natural gradient descent is an approximation to mirror descent.
@@
As the step size goes to 0, the stochasticity goes away in the online learning framework: in any time interval, as the step size goes to 0, we take more steps and so we see more new samples, effectively minimizing the population loss. Thus, instead of writing the stochastic gradient, we can write the population gradient, denoting $F$ the population loss:
@@
There are two ways to interpret what is happening here: when the step size goes to 0, both mirror descent and natural gradient descent (which are the same in the infinitesimal step size limit) converge to the Riemannian gradient flow on the population objective. (Note: this is another way to think about what we are doing with SGD, because as the step size gets smaller and smaller, we are approaching optimizing the population loss.) The other way to think about natural gradient descent and mirror descent is: we endow the space with some local geometry and take steps based on that local geometry. Natural gradient descent corresponds to a forward Euler discretization of the natural gradient flow rule (Eqn. 47):
@@ 4.1.1 Examples of Mirror Descent
which holds as long as $\Psi$ is 1-strongly convex with respect to our choice of norm. Here, $w^{\star}$ is no longer the best in the norm bounded class, since we do not wish to limit the norm explicitly. The guarantee above means that we can compete with any $w^{\star}$, but we need to pay for its complexity in the second term.
Remark
When the metric tensor doesn’t correspond to a Hessian, it is unclear wether we can identify a global complexity measure.
@@ 4.1.2 Smoothness and Batching
$\mathbb{E}\left[F(\bar{w}_{k})\right]\leq F(w^{\star})+O\left(\frac{H\Psi^{\star}(w)}{k}+\sqrt{\frac{\sigma^{2}\Psi^{\star}(w)}{k}}\right)\,,$ (53)
where $\text{where }\sigma\text{ s.t. }\mathbb{E}\left[\left|\left|\nabla f(w,z)-\nabla F(z)\right|\right|^{2}\right]\leq\sigma^{2}$
@@
We can see here is that optimization is easier than the statistical aspect. It never really helps to take more steps than the number of samples. In the non-convex case, this is substantially different: in particular, the optimization might be harder than the statistical part, and so it might be worth viewing the samples multiple times.
@@
The guarantee in Eqn. 53 holds when $H$ is the relative smoothness parameter.
@@ 4.2 General Steepest Descent
We can also think more generally about steepest descent methods that are not related to a metric,
@@ 4.3 Implicit bias of descent methods
So far, we have discussed the convex setting, and we obtained bounds combining optimization guarantees and generalization guarantees. We get generalization guarantees with respect to $w^{\star}$ considering only the dynamics of optimization. We studied the connection between the geometry of searching the space and the corresponding complexity. However, there are some limitations to what we have seen so far. First of all, this is only the convex case. Second of all, this isn’t what we’re seeing in deep learning. Here, we are not driving training error to 0. The generalization seems to rely on the fact that we are using stochasticity. That is different from what we see in the examples from the previous lectures where we train with full batch, multi-pass gradient methods but still get good generalization.
@@
$w_{\lambda}=\arg\min\langle\nabla f(w_{k},z_{k}),w\rangle+\lambda\Psi(w)\,.$
And secondly, consider $\bar{w}_{k}$, the solution achieved after $k$ iterations of mirror descent. In either of these cases, we get the same generalization guarantee, but for different solutions. In the Lipschitz case with optimally chosen $\lambda$, the distance (suboptimality, really) between $w_{\lambda}$ and $w^{\star}$ is $1/\sqrt{k}\,.$ The distance between $w^{\star}$ and $\bar{w}_{k}$ is also $1/\sqrt{k}\,,$ but the distance between $w_{\lambda},\bar{w}_{k}$ is also $1/\sqrt{k}$. We know this because we know that we can view one-pass of this as optimizing the training objective, which in $k$ steps gets suboptimality $1/\sqrt{k}\,.$ Thus, we are not saying that implicit regularization gets us to the same solution as explicit regularization, but rather we are saying that the generalization guarantees hold. If we keep repeating passes, we might get to minimizer of the training error, but it’s unclear if this is beneficial.
@@ 4.3.1 Deriving Implicit Regularization for Gradient Descent
We’ve dropped $\phi$ for ease of notation, but we should think of $x$ as being the output of some feature map. Let $w\in\mathbb{R}^{d}\,,\mbox{with}\thickspace d>>m\,.$ Let us use gradient descent:
$w_{k+1}=w_{k}-\eta_{k}\nabla\hat{F}(w_{k})\,;\quad w_{0}=0\,.$
We know from previous discussion that this will result in implicit $\ell_{2}$ regularization, but let us formally derive this. This will also allow us to understand what happens when moving to mirror descent later on.
@@
where $X\in\mathbb{R}^{m\times d}$. This simply comes from the fact that, for the model above, the gradient reads
$\nabla\hat{F}(w)$ $=\frac{1}{m}\sum_{i=1}^{m}loss^{\prime}\left(\langle w,x_{i}\rangle,y_{i}\right)x_{i}$
$=X^{\top}s,$ (55)
where we introduced the vector of residuals $s\in\mathbb{R}^{m}$ containing the entries $\frac{1}{m}loss^{\prime}\left(\langle w,x_{i}\rangle,y_{i}\right)$. This already tells us a lot about what we will get: even though we have $d$ parameters, we only will ever be in an $m$-dimensional subspace. We will also assume that we get to a global optimum. This is not always easy to show but are going to assume it. Since we are in the overparametrized setting, we know that $Xw=y\,,$ where $w$ is the predictor to which we finally converge. We claim that the fixed point of this iteration is exactly the minimizer of the following optimization problem:
@@
In general, these arguments are going to proceed by claiming that solving the first optimization problem actually implicitly solves the second optimization problem. To show this, we use the method of Karush-Kuhn-Tucker (KKT) conditions for optimality of a solution to a constrained optimization problem. The optimum for a constrained optimization problem is uniquely characterized by its KKT conditions. Let us look at the KKT conditions. We introduce dual variables $\nu$ for the constraints, giving us that the Lagrangian is:
@@
We observe that the stationarity condition shows exactly what we claimed earlier, that $w_{k}$ are in the span of the data, and the primal feasibility condition finds the interpolator. Since a point that satisfies these two conditions is optimal for the optimization problem in Eqn. 56, and since gradient descent that arrives at a 0 training error predictor satisfies these two conditions, gradient descent finds an optimal solution to the problem in Eqn. 56, i.e., a minimum Euclidean norm solution.
@@ 4.3.2 Similar Argument for Mirror Descent
With this argument in mind, we can apply a similar procedure to analyze the output of mirror descent. We are going to show that with mirror descent, we obtain a solution that implicitly minimizes the Bregman distance of the chosen potential with the starting point $w_{0}$. To show this, we will use essentially the same proof. The optimization problem is still the same as before: we are minimizing the training objective, but the optimization algorithm will differ. Now, we optimize $\hat{F}$ using mirror descent with respect to some potential function $\Psi$. Let us reproduce the iterates:
@@
$w_{k+1}=\nabla\Psi^{-1}\bigg{(}\nabla\Psi(w_{0})-\sum_{i=1}^{k}\frac{1}{\lambda_{i}}\nabla\hat{F}(w_{k})\bigg{)}$
That is, the $k+1$ iterate is the mapping back into the primal space of the accumulation of the gradients in the dual space. As before, we can see that the gradients lie in the span of the data (in the dual space). Thus, again, $w_{k}\in\mathcal{M}=\{\nabla\Psi^{-1}(\nabla\Psi(w_{0})+X^{T}s)\,\forall\,x\in\mathbb{R}^{m}\}\,.$ This is not a flat manifold (zero curvature) in the primal space, but it is in the dual space. To see what point we converge to, we use the assumption that in the end, we converge to a global optimum. This imposes $m$ linear constraints, which when intersected with an $m$-dimensional manifold gives us a unique point. To determine which unique point that is, we write an optimization problem whose KKT conditions match the two sets of constraints (global optimality and lying in the manifold):
@@
4.3.3 A General Method
We started from the optimization problem of minimizing the training objective and then saw trajectory stays within a manifold. The second set of constraints we imposed came from the global optimality of the solution reached. We then matched these two sets of constraints to KKT conditions for some other optimization problem. The goal of the next lecture will be to use this method to understand the implicit bias of optimization in various problems, and in particular attempt to highlight what parameters or architectures may lead to implicity $\ell_{2}$ regularization or $\ell_{1}$, i.e. feature learning.
@@ 5 Lecture 5: Implicit bias with linear functionals and the square loss
The main topic of this lecture is the derivation of analytical evidence to the lazy [16] and rich regimes in learning problems with gradient descent, by employing the general method described in the previous lecture. We will consider simple model for which the dynamics of gradient descent may be solved exactly and we will study these trajectories as a function of the magnitude of the initialization and model architectures, leading to various implicit biases.
@@ 5.1 Setting
We consider models parametrized by a weight vector $\mathbf{w}\in\mathbb{R}^{p}$ acting on an input space $\mathcal{X}$ and denote $F(\mathbf{w})\in\{f:\mathcal{X}\to\mathbb{R}\}$ the predictor implemented by $\mathbf{w}$, such that $F(\mathbf{w})(\mathbf{x})=f(\mathbf{w},\mathbf{x})$. We will focus on models linear in $\mathbf{x}$ represented by a linear functional in the dual space of $\mathcal{X}$, denoted $\mathcal{X}^{*}$ and represented by a vector $\boldsymbol{\beta}_{\mathbf{w}}$ such that $f(\mathbf{w},\mathbf{x})=\langle\boldsymbol{\beta}_{\mathbf{w}},\mathbf{x}\rangle$, i.e. $\boldsymbol{\beta}_{\mathbf{w}}$ is some transform, potentially non-linear of $\mathbf{w}$. We consider the supervised learning problem with $n$ sample pairs $(\mathbf{x}_{i},y_{i})$ where $y_{i}\in\mathbb{R}$ is a response vector. The parameters are learned by minimizing the empirical loss
@@
We assume the model is homogeneous of order $D$, i.e., for any constant $c>0$ $F(c\mathbf{w})=c^{D}h_{\mathbf{w}}$. The order $D$ is related to the depth of networks with homogeneous activations (e.g. a linear or Relu). A linear model is homogeneous of order $1$, a factorization model of order $2$. Our objective is three-fold :
  • first, we want to characterize the implicit bias of gradient descent in this setup. In regards to the previous lectures, is optimizing in the parameters space using gradient descent equivalent to optimizing in the functional space w.r.t. some metric tensor and potential? If this is the case, can we characterize this potential analytically?
  • secondly, is this implicit optimization problem in function space equivalent to explicit regularization in parameter space? For instance, we saw that GD on a least-squares problem converges towards the minimum $\ell_{2}$ norm solution, equivalent to explicitly penalizing the $\ell_{2}$ norm. Can we extend this picture to more general models, in particular models closer to deep learning architectures.
  • finally, we would like to study the transition between kernel (a.k.a. lazy) regime and rich feature learning regime. What parameters govern this behaviour? We have seen in the previous lectures that in many cases we obtain implicit biases that cannot be reached with kernel methods (sparsity, nuclear norm, …). Can we get a more precise picture on simple models?